OpenStudy (anonymous):

how do i simplify the square root of 81c^5,d^4

OpenStudy (anonymous):

\[\sqrt{81}=9\] \[\sqrt{c^5}=c^2\sqrt{c}\] \[\sqrt{d^4}=d^2\] mash them up

OpenStudy (anonymous):

so 9c^2square root of c d^2

OpenStudy (anonymous):

i would write as \[9c^2d^2\sqrt{c}\] here is a more modern version http://www.youtube.com/watch?v=Cw6pQUa0yS8&feature=related

OpenStudy (anonymous):

thank you and i still need some help on this stuff are you willing ?

OpenStudy (anonymous):

although i prefer the prince buster version

OpenStudy (anonymous):

suer post away

OpenStudy (anonymous):

ok so simplify the square root of 124y^6w^7

OpenStudy (anonymous):

ok step by step. first the easy one. what is \[\sqrt{y^6}\]?

OpenStudy (anonymous):

y^2?

OpenStudy (anonymous):

no. how many times does 2 go in to 6?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

that is , what is 6 divided by 2 right, 3 so \[\sqrt{y^6}=y^3\] because we know \[(y^3)^2=y^6\]

OpenStudy (anonymous):

now how about \[\sqrt{w^7}\] in this case 2 does not divide 7 evenly, but it goes in to 7 3 times with a remainder of 1, so \[\sqrt{w^7}=w^3\sqrt{w}\]

OpenStudy (anonymous):

since 2 goes in to 7 3 times, w cubed comes out, one w stays in. for another example \[\sqrt{x^{23}}=x^{11}\sqrt{x}\]

OpenStudy (anonymous):

oh got that part!

OpenStudy (anonymous):

finally \[\sqrt{124}\] now \[124=4\times 31\] so \[\sqrt{124}=\sqrt{4\times 31}=\sqrt{4}\times \sqrt{31}=2\sqrt{31}\]

OpenStudy (anonymous):

you don't need to write all the steps , that is just an explanation

OpenStudy (anonymous):

so 2square root of 31y^3w^3square root w

OpenStudy (anonymous):

put together get \[\sqrt{124y^6w^7}=2y^3w^3\sqrt{31w}\]

OpenStudy (anonymous):

put everything outside the radical first, then the radical with what is left inside

OpenStudy (anonymous):

got it thanks

OpenStudy (anonymous):

so now simplify the square root of 18/x^3

OpenStudy (anonymous):

well, this one is a little different because you have to get the \[\sqrt{x^3}\] out of the denominator

OpenStudy (anonymous):

ok and how do i do that?

OpenStudy (anonymous):

but even before we begin we can write \[\sqrt{\frac{18}{x^3}}=\frac{\sqrt{18}}{\sqrt{x^3}}\] that part is ok right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and now from before we know \[\sqrt{18}=\sqrt{9\times 2}=3\sqrt{2}\]

OpenStudy (anonymous):

similarly \[\sqrt{x^3}=x\sqrt{x}\]

OpenStudy (anonymous):

leaving us with \[\frac{3\sqrt{2}}{x\sqrt{x}}\]

OpenStudy (anonymous):

now to get the radical out of the denominator you write the following \[\frac{3\sqrt{2}}{x\sqrt{x}}=\frac{3\sqrt{2}}{x\sqrt{x}}\times \frac{\sqrt{x}}{\sqrt{x}}=\frac{3\sqrt{2x}}{x^2}\]

OpenStudy (anonymous):

now there is no radical in the denominator. hope steps are clear

OpenStudy (anonymous):

very thank you

OpenStudy (anonymous):

how do i simplify 3/5-the squaer root of 2